OpenStudy (anonymous):

What is the distance between the points (21,16) and (9,11)

OpenStudy (anonymous):

13 nm

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow Distance = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2} }\)

OpenStudy (anonymous):

13

Parth (parthkohli):

No, not 13.

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow \sqrt{(-13)^2 +(-5)^2 }}\) \(\Large \color{Black}{\Rightarrow \sqrt{169 + 25} }\) \(\Large \color{Black}{\Rightarrow \sqrt{194} }\)

OpenStudy (anonymous):

13.9